OpenStudy (lgbasallote):

a car travels along a straight line at a constant speedof 60.0 mi/h for a distance d and then another distance d in the same direction at another constant speed. the average velocity of the entire trip is 30.0 mi/h (a) what is the constant speed with which the car moved during the second distance d?

OpenStudy (lgbasallote):

i would assume that the second speed is slower than 30?

OpenStudy (lgbasallote):

and how do i set this up @rajathsbhat

OpenStudy (lgbasallote):

is it (60 + x)/2 = 30??

OpenStudy (lgbasallote):

since distance is the same

OpenStudy (anonymous):

*thinking*

OpenStudy (anonymous):

It's 20 mph.

OpenStudy (lgbasallote):

why???

OpenStudy (anonymous):

t1=d/60 t2=d/x 30=2d/t1+t2 substitute & simplify

OpenStudy (lgbasallote):

2d/ti?

OpenStudy (lgbasallote):

2d/t1?

OpenStudy (lgbasallote):

ohhh the formula...2d/t1 + t2

OpenStudy (anonymous):

yes.

OpenStudy (maheshmeghwal9):

formula is \[\color{red}{V_{av.}=\frac{2V_1V_2}{V_1+V_2}.}\]

OpenStudy (lgbasallote):

\[\large 30 = \frac{2d}{\frac{(60+x)d}{60x}}\] that right?

OpenStudy (anonymous):

yep.

OpenStudy (lgbasallote):

then \[\large 30 = \frac{60x}{60+x}\]

OpenStudy (lgbasallote):

\[180 = 30x\] xshould be 60? o.O

OpenStudy (anonymous):

1800

OpenStudy (lgbasallote):

oh wait i missed 2

OpenStudy (anonymous):

and that.

OpenStudy (lgbasallote):

\[1800 + 30x = 120x\] \[1800 = 90x\] x should be 20?

mathslover (mathslover):

the constant speed is 20 m/h. here's how you do it. it travels d distance with 60m/h, so time taken t1= d/60 it travels another d distance with a velocity say v, time taken t2= d/v hence total time taken t = t1+t2 = d/60 + d/v now avg velocity = total distance / total time = 2d/ t = 2d/ (d/60 + d/v) = 120v/ (v +60) It's given that avg velocity = 30m/h= 120v /(v+60) if you solve it, v = 20m/h i hope this is correct

OpenStudy (anonymous):

tada.

OpenStudy (lgbasallote):

great!!1 thanks

OpenStudy (anonymous):

yw :)

OpenStudy (lgbasallote):

uhmm woops forgot there was follow up question

OpenStudy (lgbasallote):

(b) suppose the second distance d were traveled in the opposite direction what is the average velocity for the trip? 0 right?

mathslover (mathslover):

here the second trip is in the opposite direction of the first trip and we're asked the vector velocity. you can close your eyes and say that the answer is zero. Because when we find the avg velocity we divide avg displacement by total time (formula in part (a) ). because the displacement is zero ( d+ (-d) = 0 ), the avg velocity of obviously zero.

mathslover (mathslover):

very right @lgbasallote gr8 work

OpenStudy (anonymous):

yes, lgba

OpenStudy (lgbasallote):

yayyy

mathslover (mathslover):

:) yahoo....

OpenStudy (lgbasallote):

(C) what is the average speed average speed is 2d/t1 + t2 right?

mathslover (mathslover):

here we have been asked the speed, so it doesn't matter in which direction we move. We have to focus on the magnitude alone. And the magnitudes have been given the same as that in part (a). hence the answer for avg speed = 30 m/h (which has been already provided in the question itself).

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

If a particle moves a distance at a speed v1 & again same distance with v2 then the formula is: -\[\color{red}{V_{av.}=\frac{2V_1V_2}{V_1+V_2}.}\]If a particle moves a distance in 2 equal intervals of time at different speeds v1 & v2 respectively then the formula is: -\[\color{red}{V_{av.}=\frac{V_1+V_2}{2}.}\]Where "Vav." is average velocity.

OpenStudy (anonymous):

wow, @mathslover how did you type so much before i could type yes?

mathslover (mathslover):

lolz .... i was prepared for that .. i had that question in my book

OpenStudy (anonymous):

hahaha..

mathslover (mathslover):

and also i am preparing for this topic ..... all lgba's question i had done just 1-2 days ago .:)

OpenStudy (lgbasallote):

so t1 = d/60 t2 = d/-20 <---because it also said that to return trip was the same constant speed as found in part (a) so then \[\frac{2d}{\frac{d}{60} - \frac{d}{20}}\] that right?

OpenStudy (anonymous):

No, for speed, there's no + or - stuff. It's + always.

mathslover (mathslover):

nope ... it is not fully right

mathslover (mathslover):

yes since speed is scalar quantity

mathslover (mathslover):

it does not include direction it only includes magnitude no direction

OpenStudy (anonymous):

t2=d/20

OpenStudy (lgbasallote):

oh so it's d/60 + d/20?

mathslover (mathslover):

right t2 = \(\frac{d}{20}\) right

OpenStudy (anonymous):

yep.

OpenStudy (maheshmeghwal9):

\[\color{green}{\text{V=speed}}\]\[\color{blue}{{\overrightarrow V}=\text{Velocity} }\]

mathslover (mathslover):

lolz @maheshmeghwal9 thankss ... i deleted that as it was taking so much space

OpenStudy (lgbasallote):

okay so summary \[\huge V_{ave} = \frac{2d}{t_1 + t_2}\] right?

OpenStudy (maheshmeghwal9):

yw:)

OpenStudy (anonymous):

yes, lgba.

mathslover (mathslover):

right

OpenStudy (lgbasallote):

wait doesnt this mean the average speed for round trip is the same as the average velocity of (a)???

OpenStudy (anonymous):

yeah, it does.

mathslover (mathslover):

yes.. it is

OpenStudy (lgbasallote):

hmm great thanks

OpenStudy (anonymous):

that's the point.Speed=don't care about the direction